OpenStudy (anonymous):

Given that CD = 15 and is a tangent to the circle, and BD = 12, which is the length of AB?

OpenStudy (raden):

where is the figure ?

OpenStudy (raden):

just use proportion : BC/AB = AB/BD 27/AB = AB/12 AB*AB = 27*12 AB^2 = 324 AB = sqrt(324) AB = 18

OpenStudy (anonymous):

thanks

OpenStudy (raden):

maybe this link can help u to understanding it http://mathworld.wolfram.com/CircleTangentLine.html

OpenStudy (raden):

welcome